# Stability-Certified On-Policy Data-Driven LQR via Recursive Learning and Policy Gradient

### v1

@@ Abstract
In this paper, we investigate a data-driven framework to solve Linear Quadratic Regulator (LQR) problems when the dynamics is unknown, with the additional challenge of providing stability certificates for the overall learning and control scheme. Specifically, in the proposed on-policy learning framework, the control input is applied to the actual (unknown) linear system while iteratively optimized. We propose a learning and control procedure, termed Relearn LQR, that combines a recursive least squares method with a direct policy search based on the gradient method. The resulting scheme is analyzed by modeling it as a feedback-interconnected nonlinear dynamical system. A Lyapunov-based approach, exploiting averaging and singular perturbations theory for nonlinear systems, allows us to provide formal stability guarantees for the whole interconnected scheme. The effectiveness of the proposed strategy is corroborated by numerical simulations, where Relearn LQR is deployed on an aircraft control problem, with both static and drifting parameters.
@@ 1 Introduction
The massive availability of data across automation and robotics applications pushed the control community to revise the traditional model-based optimal control approaches toward a learning-driven scenario. In these solutions, the control policy is iteratively updated without an explicit knowledge of the underlying dynamical system, relying solely on the collected data. Consequently, the fundamental distinction between off-policy and on-policy methods arises from the interconnection between the collected data and the current policy. More in detail, off-policy algorithms pursue a value iteration approach, and data are, in general, independent of the current policy. Conversely, on-policy algorithms employ a policy iteration framework and evaluate the performance of the current policy using data originated by the system under the same (under evaluation) policy. From the former derivations of Reinforcement Learning methods applied to Linear Quadratic (LQ) regulation [1], there has been a surge of interest within the control community toward a data-driven resolution of infinite-horizon Linear Quadratic Regulator (LQR) problem. The recent survey [2] investigates the connection between optimal control and reinforcement learning frameworks. In the context of off-policy methodologies, we find iterative methods inspired by the Kleinman algorithm [3], involving either parameter identification or direct estimate of the policy [4, 5, 6, 7, 8, 9]. The paper [5] investigates an off-policy Q-learning strategy, with an additional focus on the computational complexity. The recent works [10, 11, 12] proposes iterative algorithms that do not assume the existence of stabilizing initial policies. A model-free approach for discrete-time LQR based on reinforcement learning is studied and developed in [13]. Off-policy approaches can be further distinguished between direct, where data are used directly in the policy design phase, and indirect approaches, where a preliminary identification step is performed. Direct strategies often tackle the LQR problem by exploiting Persistently Exciting (PE) data together with semi-definite programming and Linear Matrix Inequalities (LMI) approaches, as introduced in [14]. These methodologies are thoroughly studied also in [14, 15, 16]. The work in [17] extends these concepts to unknown linear systems with switching time-varying dynamics. These LMI-based solutions also allowed for the design of control policies in case of noisy data, as explored in  [18, 19]. Direct approaches have been deployed to address also the design of robust controllers, e.g., in [20, 21]. The recent survey [22] also includes an extension to nonlinear systems. Instead, the work [23] proposes a safe-learning strategy for LQR via an indirect approach, i.e., the unknown dynamics is firstly estimated, so that the control gain is optimized on the estimated quantities. Indirect approaches are also explored in [24, 25]. Other approaches bridging the indirect and direct paradigms have been proposed in [26, 27, 28]. Another successful approach to address the LQR problem, often deployed in an off-policy setting, is represent by policy-gradient methods, see, e.g., the recent survey [29]. A complete characterization of first-order properties of the discrete-time LQR problem is given in [30]. The convergence properties of the (policy) gradient methods are thoroughly studied in [31] for discrete-time LQR. A model-free, gradient-based, strategy is proposed in [32]. While in [33], the sample complexity and convergence properties for the continuous-time case are examined. In [34] the discrete-time case is considered. Recent works also explored the non-asymptotic performances of model-free LQR algorithms. Sub-linear regret result is given in [35, 36]. Poly-logarithmic regret bounds are given in [37, 38]. The sample complexity for model-free LQR is studied in [39]. Conversely, on-policy control techniques are proposed in the continuous-time framework in [40, 41]. In [42] stability guarantees on the learning dynamics are provided. In the discrete-time context, the on-policy setting is addressed in [43] leveraging on both policy iteration and value iteration approaches. In [44], regret bounds for online LQR are provided. Importantly, although the majority of these works provide guarantees for the design of a stabilizing controller, it is still an open challenge the thorough investigation of the stability and convergence properties of the overall interconnection between the learning algorithm and the closed-loop system dynamics under the (time-varying) control policy.
@@
2 Problem Setup and Preliminaries
@@ 2.1 On-Policy Data-Driven LQR: Problem Setup
$\min_{\begin{subarray}{c}\mathrm{x}_{1},\mathrm{x}_{2},\ldots,\\ \mathrm{u}_{0},\mathrm{u}_{1},\ldots\end{subarray}}\>\>$ $\>\mathbb{E}\bigg{[}\tfrac{1}{2}\sum_{t=0}^{\infty}\Big{(}{\mathrm{\mathrm{x}}_{t}}^{\top}Q\mathrm{\mathrm{x}}_{t}+{\mathrm{u}_{t}}^{\top}R\mathrm{u}_{t}\Big{)}\bigg{]}$ (1c)
subj. to $\>\mathrm{\mathrm{x}}_{{t+1}}={A}_{\star}\mathrm{\mathrm{x}}_{t}+{B}_{\star}\mathrm{u}_{t},\qquad\qquad\mathrm{x}_{0}\sim\mathcal{X}_{0},$ (1d)
@@
The pair $({A}_{\star},{B}_{\star})$ is unknown and controllable. normal-$\blacksquare$
@@
It is worth noting that, in light of Assumption 2.1, by continuity, there exists $\Theta\subset\mathbb{R}^{(n+m)\times n}$ such that $[{A}_{\star}\hskip 1.42271pt{B}_{\star}]^{\top}\in\Theta$ and $(A,B)$ is controllable for all $[A\hskip 1.42271ptB]^{\top}\in\Theta$. We denote by $\Theta_{b}\subseteq\Theta$ the largest ball contained in $\Theta$ and centered in $\theta^{\star}$. It is well-known that, when $({A}_{\star},{B}_{\star})$ are known the optimal solution to problem (2.1) is given by a linear time-invariant policy $\mathrm{u}_{t}=K^{\star}\mathrm{\mathrm{x}}_{t}$ with $K^{\star}\in\mathbb{R}^{{m}\times{n}}$ given by
@@
where $P^{\star}\in\mathbb{R}^{{n}\times{n}}$ solves the Discrete-time Algebraic Riccati Equation associated to problem (2.1), see [45].
In this paper, we are interested in devising a data-driven on-policy strategy to design a state-feedback controller solution of (2.1).
Hence, the problem we address is the following.
Stability-certified on-policy LQR: design a learning and control scheme capable of
@@
    learning the optimal policy solution of problem (2.1);
@@
    actuating the (real) system with the currently available state-feedback policy;
@@
    ensuring asymptotic stability of the closed-loop learning and control system.
The problem setup is illustrated in Figure 1, where $\pi_{t}(\mathrm{\mathrm{x}}_{t})$ represents the policy actuated on the real system.
@@ 2.2 Preliminaries: Model-based Gradient Method for LQR
Next, we recall the key ingredients for devising a model-based gradient method to address problem (2.1).
@@ 2.2.1 Model-based reduced problem formulation
First of all, we recall an equivalent (unconstrained) formulation of problem (2.1) that explicitly imposes the linear feedback structure to the optimal input and is amenable for gradient-based algorithmic solutions. Problem (2.1) is rewritten by substituting in the dynamics and in the cost function the input in linear feedback form $\mathrm{u}_{t}=K\mathrm{\mathrm{x}}_{t}$, where $K\in\mathbb{R}^{{m}\times{n}}$ is to be computed. First of all, given any gain $K$, the original (open-loop) dynamics (1d) admits the closed-loop formulation $\mathrm{\mathrm{x}}_{{t+1}}=({A}_{\star}+{B}_{\star}K)\mathrm{\mathrm{x}}_{t}$. So that, for all $t\geq 0$, the state is uniquely determined as
@@
Hence, leveraging on (4), assuming that $\mathcal{X}_{0}$ is a uniform distribution on the unit sphere, and taking the expected value on the initial condition, problem (2.1) can be rewritten as
@@
$J(K,\theta^{\star})=\tfrac{1}{2}\operatorname{Tr}\left(\sum_{t=0}^{\infty}({A}_{\star}+{B}_{\star}K)^{t,\top}(Q+K^{\top}RK)({A}_{\star}+{B}_{\star}K)^{t}\right).$ (6)

### v2

@@ Abstract
In this paper, we investigate a data-driven framework to solve Linear Quadratic Regulator (LQR) problems when the dynamics is unknown, with the additional challenge of providing stability certificates for the overall learning and control scheme. Specifically, in the proposed on-policy learning framework, the control input is applied to the actual (unknown) linear system while iteratively optimized. We propose a learning and control procedure, termed Relearn LQR, that combines a recursive least squares method with a direct policy search based on the gradient method. The resulting scheme is analyzed by modeling it as a feedback-interconnected nonlinear dynamical system. A Lyapunov-based approach, exploiting averaging and timescale separation theories for nonlinear systems, allows us to provide formal stability guarantees for the whole interconnected scheme. The effectiveness of the proposed strategy is corroborated by numerical simulations, where Relearn LQR is deployed on an aircraft control problem, with both static and drifting parameters.
@@ 1 Introduction
The massive availability of data across automation and robotics applications pushed the control community to revise the traditional model-based optimal control approaches toward a learning-driven scenario. In these solutions, the control policy is iteratively updated without an explicit knowledge of the underlying dynamical system, relying solely on the collected data. Consequently, the fundamental distinction between off-policy and on-policy methods arises from the interconnection between the collected data and the current policy. More in detail, off-policy algorithms pursue a value iteration approach, and data are, in general, independent of the current policy. Conversely, on-policy algorithms employ a policy iteration framework and evaluate the performance of the current policy using data originated by the system under the same (under evaluation) policy. From the former derivations of Reinforcement Learning methods applied to Linear Quadratic (LQ) regulation [1], there has been a surge of interest within the control community toward a data-driven resolution of infinite-horizon Linear Quadratic Regulator (LQR) problem. The recent survey [2] investigates the connection between optimal control and reinforcement learning frameworks.
@@
In the context of off-policy methodologies, we find iterative methods inspired by the Kleinman algorithm [3], involving either parameter identification or direct estimate of the policy [4, 5, 6, 7, 8, 9]. The paper [5] investigates an off-policy Q-learning strategy, with an additional focus on the computational complexity. The recent works [10, 11, 12] proposes iterative algorithms that do not assume the existence of stabilizing initial policies. A model-free approach for discrete-time LQR based on reinforcement learning is studied and developed in [13]. Off-policy approaches can be further distinguished between direct, where data are used directly in the policy design phase, and indirect approaches, where a preliminary identification step is performed. Direct strategies often tackle the LQR problem by exploiting Persistently Exciting (PE) data together with semi-definite programming and Linear Matrix Inequalities (LMI) approaches, as introduced in [14]. These methodologies are thoroughly studied also in [14, 15, 16]. The work in [17] extends these concepts to unknown linear systems with switching time-varying dynamics. These LMI-based solutions also allowed for the design of control policies in case of noisy data, as explored in  [18, 19]. Direct approaches have been deployed to address also the design of robust controllers, e.g., in [20, 21]. The recent survey [22] also includes an extension to nonlinear systems. Instead, the work [23] proposes a safe-learning strategy for LQR via an indirect approach, i.e., the unknown dynamics is firstly estimated, so that the control gain is optimized on the estimated quantities. Indirect approaches are also explored in [24, 25]. Other approaches bridging the indirect and direct paradigms have been proposed in [26, 27, 28]. Another successful approach to address the LQR problem, often deployed in an off-policy setting, is represented by policy-gradient methods, see, e.g., the recent survey [29]. A complete characterization of first-order properties of the discrete-time LQR problem is given in [30]. The convergence properties of the (policy) gradient methods are thoroughly studied in [31] for discrete-time LQR. A model-free, gradient-based, strategy is proposed in [32]. While in [33], the sample complexity and convergence properties for the continuous-time case are examined. In [34] the discrete-time case is considered. Recent works also explored the non-asymptotic performances of model-free LQR algorithms. Sub-linear regret result is given in [35, 36]. Poly-logarithmic regret bounds are given in [37, 38]. The sample complexity for model-free LQR is studied in [39].
Conversely, on-policy control techniques are proposed in the continuous-time framework in [40, 41]. In [42] stability guarantees on the learning dynamics are provided. In the discrete-time context, the on-policy setting is addressed in [43] leveraging on both policy iteration and value iteration approaches. In [44], regret bounds for online LQR are provided.
We underline that while the mentioned works offer guarantees for (asymptotically) obtaining stabilizing (possibly non-optimal) controllers, a thorough investigation into the stability and convergence properties of the interlacing of optimization, learning and control, governed by a time-varying and nonlinear dynamics, remains an open challenge.
@@
The key strength of the proposed architecture (and analysis) lies in its inherent robustness. The stability properties we prove ensure that the algorithm remains effective even in non-nominal conditions, where system and cost matrices change over time and/or disturbances affect the plant and the measurements. In such scenarios, the closed-loop system adapts dynamically, restoring optimality without requiring any restart of either the optimization or the learning process.
To the best of the authors’ knowledge, our attempt is the first one (in discrete-time LQR) that provides a stability certificate for the overall closed-loop system that simultaneously addresses optimization, learning, and control tasks. More in detail, existing data-driven approaches (see, e.g., [14]) are not on-policy, namely, they are characterized by two distinct phases in which system samples are collected and then used to find the optimal gain. An alternative, popular approach (see, e.g., [31, 7]) is based on improving the tentative policy by performing so-called experiments of the actuated system to evaluate it. The main drawback of these policy-iteration approaches is that they are not online, namely, the real plant needs to be repeatedly initialized and actuated for a number of samples at each algorithm iteration. Another branch of literature relies on the certainty equivalence principle, see, e.g., [24], which relates model and gain matrices accuracy and, thus, treats the optimization task as a mere corollary of the learning phase. However, to the best of the authors’ knowledge, also this class of works do not focus on stability aspects of the overall closed-loop dynamics that tries to concurrently optimize the gain, learn the system matrices, and control the real plant.
@@
2 Preliminaries and Problem Setup
In this section we present some useful preliminaries and describe the problem we aim at investigating.
2.1 Preliminaries on averaging theory for two-time-scale systems
We report [45, Theorem 2.2.4], which is a useful result in the context of averaging theory for two-time-scale systems. Consider the time-varying system
$\chi_{t+1}$ $=\mathcal{A}(\mathrm{z}_{t})\chi_{t}+h(\mathrm{z}_{t},t)+\epsilon g(\chi_{t},\mathrm{z}_{t},t)$ (1a)
$\mathrm{z}_{{t+1}}$ $=\mathrm{z}_{t}+\epsilon f(\chi_{t},\mathrm{z}_{t},t),$ (1b)
with $\chi_{t}\in\mathbb{R}^{n_{\chi}}$, $\mathrm{z}_{t}\in\mathbb{R}^{n_{\mathrm{z}}}$, $g:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, $f:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, and $\mathcal{A}:\mathbb{R}^{n_{\mathrm{z}}}\to\mathbb{R}^{n_{\chi}\times n}$. Moreover, $\epsilon>0$ is a tuning parameter that is useful to arbitrarily reduce the variations over time of subsystem (1b), which is therefore typically referred to as the slow subsystem, with $\mathrm{z}_{t}$ being the slow state. Coherently, subsystem (1a) is referred to as the fast subsystem, with $\chi_{t}$ being the fast state. Notice that the analysis also relies on the investigation of a time-invariant auxiliary system associated to the slow dynamics, termed the averaged system.
In the following, we report some conditions needed to establish the stability results of the interconnected time-varying system (1) that will be used in the forthcoming algorithmic analysis.
First, we need the following regularity conditions on the vector fields of system (1).
Assumption 2.1
There exists $r$ such that $f$, $g$, and $h$ are Lipschitz continuous into $\mathcal{B}_{r}(0_{n_{\chi}+n_{\mathrm{z}}})$. $\blacksquare$
Second, we assume that the origin is an equilibrium point for system (1).
Assumption 2.2
It holds $h(0,t)=0$, $g(0,0,t)=0$, and $f(0,0,t)=0$ for all $t\in\mathbb{N}$. $\blacksquare$
Third, we characterize the matrix function $\mathcal{A}(\mathrm{z})$ of (1a).
Assumption 2.3
Moreover, $\mathcal{A}$ is differentiable and there exists $k_{a}>0$ such that
for all $i\in\{1,\ldots,m\}$ and $\mathrm{z}\in\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$. $\blacksquare$
Fourth, we impose that the vector field characterizing the averaged system, say $f^{\textsc{av}}:\mathbb{R}^{n}_{\mathrm{z}}\to\mathbb{R}^{n}_{\mathrm{z}}$, is well-posed.
Assumption 2.4
The function $f$ is piecewise continuous in $t$ and the following limit
$f^{\textsc{av}}(\mathrm{z}):=\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(0,\mathrm{z},\tau)$ (2)
exists uniformly in $\bar{t}\in\mathbb{N}$ and for all $\mathrm{z}\in\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$. $\blacksquare$
Finally, we impose the following conditions on the difference between the time-varying vector field $f(\cdot,\cdot,t)$ and the averaged one $f^{\textsc{av}}$.
Assumption 2.5
Consider $f^{\textsc{av}}$ as defined in (2) and let $\Delta f:\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\mathrm{z}}}$ be defined as
$\Delta f(\mathrm{z},t):=f(0,\mathrm{z},t)-f^{\textsc{av}}(\mathrm{z}).$
$\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta f}(\mathrm{z},\tau)\right\|$ $\leq\nu(T)\left\|\mathrm{z}\right\|$
$\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta f}(\mathrm{z},\tau)}{\partial\mathrm{z}}\right\|$ $\leq\nu(T),$
uniformly in $\bar{t}\in\mathbb{N}$ and for all $\mathrm{z}\in\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$. $\blacksquare$
We are now ready to introduce the mentioned averaged system associated to (1). It reads as
$\tilde{\mathrm{z}}^{\textsc{av}}_{t+1}=\tilde{\mathrm{z}}^{\textsc{av}}_{t}+\epsilon f^{\textsc{av}}(\tilde{\mathrm{z}}^{\textsc{av}}_{t}).$ (4)
Notice that the averaged system effectively “neglects” the time variability of $f(\cdot,\cdot,t)$ and assumes that the fast state $\chi_{t}$ is at the origin. The underlying idea is that, as $\epsilon$ becomes smaller, a timescale separation can be established between the evolution of $\mathrm{z}_{t}$ and the variations of $f(\cdot,\cdot,t)$ and $\chi_{t}$. Operatively, for sufficiently small $\epsilon$, the following result claims that exponential stability of the origin for (4) is enough to have exponential stability of the origin for the interconnected time-varying system (1).
Theorem 2.6
[45, Theorem 2.2.4] Consider system (1) and let Assumptions 2.12.22.32.4 and 2.5 hold. If there exists $\epsilon_{0}>0$ such that, for all $\epsilon\in(0,\epsilon_{0})$, the origin is exponentially stable for system (4), then there exists $\epsilon_{1}>0$ such that, for all $\epsilon\in(0,\epsilon_{1})$, the origin is an exponentially stable equilibrium of system (1). $\blacksquare$
@@ 2.2 On-Policy Data-Driven LQR: Problem Setup
$\min_{\begin{subarray}{c}\mathrm{x}_{1},\mathrm{x}_{2},\ldots,\\ \mathrm{u}_{0},\mathrm{u}_{1},\ldots\end{subarray}}\>\>$ $\>\mathbb{E}\bigg{[}\tfrac{1}{2}\sum_{t=0}^{\infty}\Big{(}{\mathrm{\mathrm{x}}_{t}}^{\top}Q\mathrm{\mathrm{x}}_{t}+{\mathrm{u}_{t}}^{\top}R\mathrm{u}_{t}\Big{)}\bigg{]}$ (5a)
subj. to $\>\mathrm{\mathrm{x}}_{{t+1}}={A}_{\star}\mathrm{\mathrm{x}}_{t}+{B}_{\star}\mathrm{u}_{t},\qquad\qquad\mathrm{x}_{0}\sim\mathcal{X}_{0},$ (5b)
@@
The pair $({A}_{\star},{B}_{\star})$ is unknown and controllable. $\blacksquare$
Notice that the positive definiteness of $Q$ can be relaxed to mere positive semidefiniteness, provided that the pair $({A}_{\star},Q_{0})$ is detectable with $Q:=Q_{0}^{\top}Q_{0}$.
@@
It is worth noting that, in light of Assumption 2.7, by continuity, there exists $\Theta\subset\mathbb{R}^{(n+m)\times n}$ such that $[{A}_{\star}\hskip 1.42271pt{B}_{\star}]^{\top}\in\Theta$ and $(A,B)$ is controllable for all $[A\hskip 1.42271ptB]^{\top}\in\Theta$. We denote by $\Theta_{b}\subseteq\Theta$ the largest ball contained in $\Theta$ and centered in $\theta^{\star}$. It is well-known that, when $({A}_{\star},{B}_{\star})$ are known the optimal solution to problem (5) is given by a linear time-invariant policy $\mathrm{u}_{t}=K^{\star}\mathrm{\mathrm{x}}_{t}$ with $K^{\star}\in\mathbb{R}^{{m}\times{n}}$ given by
@@
where $P^{\star}\in\mathbb{R}^{{n}\times{n}}$ solves the Discrete-time Algebraic Riccati Equation associated to problem (5), see [46].
In this paper, we are interested in devising a data-driven on-policy strategy to design a state-feedback controller solution of (5). Hence, the problem we address can be posed as follows.
Stability-certified on-policy LQR: design a learning and control scheme that is capable of
@@
    learning the optimal policy solution to problem (5),
@@
    estimating the unknown system matrices,
@@
    actuating the (real) system with the currently available state-feedback policy,
while ensuring asymptotic stability of the closed-loop learning and control system.
As shown in Figure 1, the proposed scheme aims to iteratively compute a tentative policy, say $\pi_{t}(\cdot)$, to be actuated on the real, though unknown, system. One single sample of the system evolution is then collected and fed into a learning mechanism that refines the system matrices, which, in turn, are used to improve the quality of the previously actuated policy. The distinctive feature of the proposed approach is that the aforementioned steps are interwoven rather than temporally separate, as is typically done in the literature.
@@ 2.3 Preliminaries: Model-based Gradient Method for LQR
Next, we recall the key ingredients for devising a model-based gradient method to address problem (5).
@@ 2.3.1 Model-based reduced problem formulation
First of all, we recall an equivalent (unconstrained) formulation of problem (5) that explicitly imposes the linear feedback structure to the optimal input and is amenable for gradient-based algorithmic solutions. Problem (5) is rewritten by substituting in the dynamics and in the cost function the input in linear feedback form $\mathrm{u}_{t}=K\mathrm{\mathrm{x}}_{t}$, where $K\in\mathbb{R}^{{m}\times{n}}$ is to be computed. First of all, given any gain $K$, the original (open-loop) dynamics (5b) admits the closed-loop formulation $\mathrm{\mathrm{x}}_{{t+1}}=({A}_{\star}+{B}_{\star}K)\mathrm{\mathrm{x}}_{t}$. So that, for all $t\geq 0$, the state is uniquely determined as
@@
Hence, leveraging on (7), assuming, without loss of generality, that $\mathcal{X}_{0}$ is a uniform distribution on the unit sphere, and taking the expected value on the initial condition, problem (5) can be rewritten as
@@
$J(K,\theta^{\star})=$ (9)
$\tfrac{1}{2}\operatorname{Tr}\left(\sum_{t=0}^{\infty}({A}_{\star}+{B}_{\star}K)^{t,\top}(Q+K^{\top}RK)({A}_{\star}+{B}_{\star}K)^{t}\right).$
@@
Remark 2.8
Due to the linearity of the expected value operator and the properties of the trace, the formulation in (9) holds up to a constant scaling factor for any probability distribution $\mathcal{X}_{0}$ with a well-defined second moment.